Anika is making nut mixture to sell at a local farmers market. She mixes 4 pounds of peanuts with a nut mixture that is 55% peanuts. The resulting mixture is 70% peanuts .how many pounds of nut mixture does she make?
Answer by ankor@dixie-net.com(22740)

Anika is making nut mixture to sell at a local farmers market.
She mixes 4 pounds of peanuts with a nut mixture that is 55% peanuts.
The resulting mixture is 70% peanuts. how many pounds of nut mixture does she make?
:
let x = amt of 55% peanuts
then
(x+4) = amt of resulting mixture
:
the decimal equiv of the per cent has to be used here
4 + .55x = .70(x+4)
4 + .55x = .70x + 2.8
4 - 2.8 = .70x - .55x
1.2 = .15x
x = 1.2/.15
x = 8 lb of 55% nuts required
then
4 + 8 = 12 lbs, resulting nut mixture made